ION ORGANIZATION AND REVERSED ELECTRIC FIELD AT AIR/AQUEOUS INTERFACES RE-VEALED BY HETERODYNE-DETECTED SUM FREQUENCY GENERATION SPECTROSCOPY

WEI HUA, ZISHUAI HUANG, AARON M. JUBB, \underline{HEATHER C. ALLEN}, Department of Chemistry, The Ohio State University, 100 W. 18th Ave., Columbus, OH, 43210.

Sum frequency generation (SFG) is a second order optical spectroscopy that probes regions of non centrosymmetry, interfaces, and allows for the understanding of molecular organization at air/aqueous interfaces. An overview of our work in this area is presented with emphasis on phase-sensitive SFG (PS-SFG) spectroscopy. PS-SFG is a variant of SFG and is used in our laboratory to investigate the average direction of the transition dipole of interfacial water molecules. The orientation of water at air/aqueous inorganic salts interfaces of CaCl_{2}, NaCl, Na_{2}SO_{4}, (NH_{4})_{2}SO_{4}, and Na_{2}CO_{3} is inferred from the direct measurement of the transition dipole moment. We find that charge separation at the air/water interface is most obvious for the aqueous ammonium sulfate solution where the local electric field has a greater magnitude at this interface relative to the other salt solutions. The magnitude of the electric field in the surface extending to the subsurface regions decreases in the order: (NH_{4})_{2}SO_{4} > Na_{2}SO_{4} > Na_{2}CO_{3} ≥ CaCl_{2} > NaCl; the electric field is opposite in direction for the sulfates and carbonate relative to the chloride salts.